OpenStudy (anonymous):

6 square root of x plus 1 times 5 square root of x - 2

OpenStudy (anonymous):

is that \[\huge 6\sqrt{x+1} \;\;\times\;\; 5\sqrt{x-2}\]

OpenStudy (anonymous):

if the above is correct you can factor away the five and the 6 and multiply them bringing the equation to its simplest form at 30sqrt(x+1)sqrt(x-2)